what is the formula for sin law and cosin law and mention it fast

Grade:11

1 Answers

SHAIK HAFEEZUL KAREEM
8 years ago
sin law=sinaAa=sinB/b=sinC/c...................and cos law=c2=a2+b2+cosC these are the cosine and sine laws